You buy a bond with a $1,000 par value today for a price of $835. The bond has 6 years to maturity and makes annual coupon payments of $67 per year. You hold the bond to maturity, but you do not reinvest any of your coupons. What was your effective EAR over the holding period?

7.68%

11.19%

9.02%

10.55%
Answer #1

The amount of coupons received = 6*67 = 402
Number of years = 6
Price today = 835
Return over 6 years = (1000-835+402)/835 = 67.904%

Return over 1 year = (1+67.904%)^(1/6) -1 = 9.02%
(Option c)

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well